# V2X-Radar: A Multi-modal Dataset with 4D Radar for Cooperative Perception

### v2

@@ 6 Conclusion
A Appendix
This supplementary document is organized as follows:
  • We offer further details on the time synchronization of sensors on vehicle and roadside platforms in Appendix B.
  • We provide supplementary benchmark results, which include roadside 3D object detection on V2X-Radar-I sub-dataset and vehicle-side 3D object detection on V2X-Radar-V sub-dataset, both evaluated under heterogeneous splits in Appendix C.
  • We provide visualizations of different data collection scenarios, including corner cases for single-vehicle autonomous driving and a variety of time-of-day and weather conditions in Appendix D.
B Synchronization Details
Time synchronization enables the simultaneous sampling of data from the same scene by various sensors on both the vehicle platform and roadside units. For a collaborative perception dataset, achieving precise time alignment is critical—not only among the diverse sensors within a single platform (be it the vehicle or the roadside unit) but also across sensors spanning both platforms.
C More Benchmark Results
We present benchmark results for single-agent 3D object detection, which include roadside 3D object detection on the V2X-Radar-I sub-dataset and single-vehicle 3D object detection on the V2X-Radar-V sub-dataset. All methods are evaluated using the heterogeneous split.
C.1 Roadside 3D Object Detection
Tab. 6 presents a quantitative evaluation of multiple roadside 3D object detection models, leveraging our V2X-Radar-I sub-dataset under a heterogeneous split configuration. The results reveal that:
  • LiDAR-based methods achieve the best performance, while 4D Radar and camera-based methods are less effective than LiDAR. Furthermore, the performance of 4D Radar and camera-based methods is comparable; 4D Radar excels at IoU=0.5, while camera-based methods perform better at the stricter IoU=0.7.
  • Compared to the results in Tab.4, both camera-based and 4D Radar-based roadside methods under the heterogeneous split show a significant decline in performance, highlighting that scenario generalization presents a challenge for both methods.
C.2 Single-vehicle 3D Object Detection
Tab. 7 presents comparative experiments of various single-vehicle 3D object detection models utilizing our V2X-Radar-V sub-dataset under a heterogeneous split. It is evident from the table that:
  • LiDAR-based methods demonstrate the highest performance. 4D Radar-based methods outperform camera-based methods but do not match the effectiveness of LiDAR-based approaches, exhibiting slight deviations from the results in Tab. 6.
  • In comparison to the findings presented in Tab. 5, both camera-based and 4D Radar single-vehicle approaches under the heterogeneous split exhibit markedly reduced performance. This underscores the substantial difficulties these methods face in achieving robust generalization across diverse scenarios.
D Data Collection Scenarios
Our data collection encompasses various weather conditions and times of day, focusing on challenging corner cases for single-vehicle autonomous driving. Fig. 8 presents a visual depiction of various challenging corner cases, such as occluded vehicles at a public intersection, partially hidden cyclists at a public T-junction, concealed pedestrians within a restricted park environment, and obscured cyclists at a campus intersection. Fig. 9 illustrates the scenarios employed for data acquisition, spanning diverse times of day and meteorological conditions. As highlighted in Fig. 9(a), the temporal segments include morning daytime, afternoon daytime, dusk, and nighttime. As shown in Figure 9(b), the weather conditions range from clear sunny skies to fog, rain, and snow.

### v3

@@ Appendix F Data Collection Scenarios
Figure 10 provides a comprehensive visualization of representative corner cases captured in the V2X-Radar, reflecting the diverse and complex challenges encountered in real-world autonomous driving. These include occluded vehicles at public intersections, partially hidden cyclists at T-junctions, concealed pedestrians within restricted park areas, and obscured cyclists at campus crossings.
Figure 10: Diverse corner cases for single-vehicle autonomous driving. Each case is depicted by roadside multi-view images on the left and the vehicle-side image on the right. We use a red circle to mark the occluded objects. The red circle highlights critical objects that are occluded from the vehicle-mounted perspective but remain observable from the roadside viewpoint.